OpenStudy (anonymous):

What is the simplified form of 9x^2-25/3x-5?

OpenStudy (anonymous):

3x + 5

OpenStudy (saifoo.khan):

\[\frac{(3x-5)(3x+5)}{3x-5}\]

OpenStudy (saifoo.khan):

both 3x - 5 cancels out, finally.

OpenStudy (saifoo.khan):

\[3x + 5\]

OpenStudy (anonymous):

\[\frac{\cancel{(3x-5)}(3x+5)}{\cancel{3x-5}}\]

OpenStudy (saifoo.khan):

Woow! that is cool.

OpenStudy (anonymous):

ho ho ho ho an new trick!

OpenStudy (anonymous):

but i absolutely hate it when math teachers do that. any one should be able to read. you certainly wouldn't do it in an english paper would you?

OpenStudy (saifoo.khan):

\[\Huge \frac{\cancel{new-trick}}{\cancel{new-trick}} \]